Use the following information to answer exercise 1 questions:

A population-based cohort study was conducted in an outlying area of Brisbane. The findings included that:

  • 45% of the population is overweight, and
  • 72.5% of the people with adult-onset diabetes were overweight;
  • The incidence of adult-onset diabetes was 370 per 100,000 person-years in the obese, 40 per 100,000 person-years in non-obese and 105 per 100,000 person-years in the whole population.
    1. Calculate the RR:

  1. Calculate the Risk Difference and describe your findings

  1. What other terms are used to describe Risk Difference?
Answer #1

Ans 1:

Relative risk = incidence of adult-onset diabetes in Obese to non-obese = 370/40 = 9.25

Ans 2:

Risk difference = 370 - 40 = 330 per 100,000 person years.

This is called a perspective risk difference as it does not tell the differences in percentages of people with risks

Ans 3:

Absolute difference, odds ratio are other terms that can be used to describe the absolute or relative differences in risk